Use Mathematical Induction to prove the Binomial Theorem.
1
Expert's answer
2020-11-17T17:05:26-0500

Let us give a proof of the Binomial Theorem using mathematical induction. We will need to use Pascal's identity in the form:


"\left(\begin{array}{c}n\\r-1\end{array}\right)+\left(\begin{array}{c}n\\r\end{array}\right)=\left(\begin{array}{c}n+1\\r\end{array}\right)" for "0<r\le n".


We aim to prove that


"(a+b)^n=a^n+\left(\begin{array}{c}n\\1\end{array}\right)a^{n−1}b+\left(\begin{array}{c}n\\2\end{array}\right)a^{n−2}b^2+⋯+\left(\begin{array}{c}n\\r\end{array}\right)a^{n−r}b^r +⋯+\left(\begin{array}{c}n\\n-1\end{array}\right)ab^{n−1}+ b^n."


We first note that the result is true for n=1 and n=2: "(a+b)^1=a+b" and "(a+b)^2=a^2+2ab+b^2=a^2+\left(\begin{array}{c}2\\1\end{array}\right)ab+b^2".


Let "k" be a positive integer with "k\u22652" for which the statement is true. So


"(a+b)^k=a^k+\left(\begin{array}{c}k\\1\end{array}\right)a^{k−1}b+\left(\begin{array}{c}k\\2\end{array}\right)a^{k−2}b^2+⋯+\left(\begin{array}{c}k\\r\end{array}\right)a^{k−r}b^r +⋯+\left(\begin{array}{c}k\\k-1\end{array}\right)ab^{k−1}+ b^k"


Now consider the expansion


"(a+b)^{k+1}=(a+b)(a+b)^k="


"=(a+b)(a^k+\left(\begin{array}{c}k\\1\end{array}\right)a^{k−1}b+\left(\begin{array}{c}k\\2\end{array}\right)a^{k−2}b^2+⋯+\left(\begin{array}{c}k\\r\end{array}\right)a^{k−r}b^r +⋯+\left(\begin{array}{c}k\\k-1\end{array}\right)ab^{k−1}+ b^k)="


"=a^{k+1}+\left[1+\left(\begin{array}{c}k\\1\end{array}\right)\right]a^{k}b+\left[\left(\begin{array}{c}k\\1\end{array}\right)+\left(\begin{array}{c}k\\2\end{array}\right)\right]a^{k−1}b^2+⋯+\left[\left(\begin{array}{c}k\\r-1\end{array}\right)+\left(\begin{array}{c}k\\r\end{array}\right)\right]a^{k+1−r}b^r +⋯+\left[\left(\begin{array}{c}k\\k-1\end{array}\right)+1\right]ab^{k}+ b^{k+1}"


From Pascal's identity, it follows that


"(a+b)^{k+1}=a^{k+1}+\left(\begin{array}{c}k+1\\1\end{array}\right)a^{k}b+\left(\begin{array}{c}k+1\\2\end{array}\right)a^{k−1}b^2+⋯+\left(\begin{array}{c}k+1\\r\end{array}\right)a^{k+1−r}b^r +⋯+\left(\begin{array}{c}k+1\\k\end{array}\right)ab^{k}+ b^{k+1}"


Hence the result is true for "k+1". By induction, the result is true for all positive integers "n".